### v1

# Context-Aware Deep Learning for Robust Channel Extrapolation in Fluid Antenna Systems

Abstract
Fluid antenna systems (FAS) offer remarkable spatial flexibility but face significant challenges in acquiring high-resolution channel state information (CSI), leading to considerable overhead. To address this issue, we propose CANet, a robust deep learning model for channel extrapolation in FAS. CANet combines context-adaptive modeling with a cross-scale attention mechanism and is built on a ConvNeXt v2 backbone to improve extrapolation accuracy for unobserved antenna ports. To further enhance robustness, we introduce a novel spatial amplitude perturbation strategy, inspired by frequency-domain augmentation techniques in image processing. This motivates the incorporation of a Fourier-domain loss function, capturing frequency-domain consistency, alongside a spectral structure consistency loss that reinforces learning stability under perturbations. Our simulation results demonstrate that CANet outperforms benchmark models across a wide range of signal-to-noise ratio (SNR) levels.
Index Terms:
Fluid antenna system (FAS), channel extrapolation, deep learning, amplitude perturbation.
I Introduction
With the advent of sixth-generation (6G) mobile communications, the demand for higher capacity, wider coverage, and greater flexibility in wireless systems is accelerating. Meeting these ambitious requirements necessitates the development of advanced physical-layer technologies. Among them, the fluid antenna system (FAS) [1, 2] has emerged as a promising antenna architecture that introduces new spatial degrees of freedom (DoF) into wireless communication. FAS represents the capability of shape-and-position-reconfigurable antenna technology to empower the physical layer. By leveraging mechanical motion [3], liquid metals [4], reconfigurable pixel-based structures [5] or metamaterials [6], FAS enables dynamic port selection from a massive number of predefined locations. This adaptability facilitates spatial diversity gains in nearly continuous space and enhances the system’s capability to respond intelligently to the wireless channel. In [7], Lu et al. explained FAS from the viewpoint of electromagnetics.
Since the introduction of FAS by Wong et al. in [8, 9], there have been lots of efforts in understanding the achievable performance of FAS [10, 11]. The use of FAS has also been considered for secrecy communications [12], reconfigurable intelligent surface (RIS) systems [13], index modulation [14, 15], wideband communication [16], and integrated sensing and communication (ISAC) [17]. Capacity maximization for FAS channels has also been studied using statistical channel state information (CSI) [18] and instantaneous CSI [19]. The diversity-multiplexing trade-off for multiple-input multiple-output (MIMO) FAS channels was analyzed in [20].
Nonetheless, the promising performance of FAS relies on the availability of CSI and the high-resolution nature of FAS also leads to substantial overhead in acquiring CSI, including pilot overhead, feedback latency, and energy consumption. In recent years, several studies explored the integration of physical modeling and data-driven approaches for CSI prediction [21, 22]. For instance, [23] proposed an inference algorithm to enable high-accuracy extrapolation in a massive FAS using only a small number of observed ports. Additionally, in [24], an extrapolation scheme based on masked language models, which reconstructs the full channel from a limited number of observed port CSI, was proposed. Going one step further, the approach in [25] integrates the diffusion mechanism of graph neural networks with an asymmetric masked autoencoder architecture, significantly improving both extrapolation accuracy and flexibility in high-resolution FAS settings. Prior to these, there were attempts that exploited channel sparsity [26, 27] and used Bayesian approaches [28, 29]. CSI prediction for multiuser FAS using deep learning was also reported in [30]. Overall, however, reliable CSI acquisition seems only possible at high signal-to-noise ratio (SNR) and it is important to study enhancement techniques that work well in noisy situations.
Motivated by the above, in this correspondence, we propose CANet, a novel deep learning framework designed for channel extrapolation in FAS. CANet adaptively infers the CSI of the unobserved ports with varying positions within a FAS, and maintains robust performance across different SNR scenarios. The architecture of CANet incorporates a context adaptive block (CAB), a cross-scale contextual attention (CSCA) mechanism, and a ConvNeXt v2 backbone enhanced with dropout regularization. This combination empowers CANet to process unobserved port positions both locally and globally, significantly improving the extraction of CSI features. To enhance model robustness further, we also introduce a spatial amplitude perturbation strategy. Besides, we refine the loss function by incorporating a fast Fourier transform (FFT) loss, which collaborates with a spectral structure consistency loss to reinforce model stability. Numerical results indicate that CANet surpasses benchmark models, demonstrating superior channel extrapolation robustness across varying noise levels.
II Channel Extrapolation for FAS
Consider a downlink system where a base station (BS) with $M_{t}$ fixed-position antennas (FPAs) communicates with $U_{t}$ user terminals on the same time-frequency resource unit. Each user terminal employs a 2D FAS with $N_{s}=N_{x}\times N_{y}$ switchable positions (or ports), which are evenly distributed over a planar surface of physical dimensions $W_{s}=W_{x}\times W_{y}$. In a 3D rich scattering environment, the spatial correlation between port $(n_{x}^{s},n_{y}^{s})$ and port $(\tilde{n}_{x}^{s},\tilde{n}_{y}^{s})$ is given by
$\begin{array}[]{l}J_{\left(n_{x},n_{y}\right),\left(\tilde{n}_{x},\tilde{n}_{y}\right)}\\ \quad=j_{0}\left(\frac{2\pi}{\lambda}\sqrt{\left(\frac{\left|n_{x}-\tilde{n}_{x}\right|}{N_{x}-1}W_{x}\right)^{2}+\left(\frac{\left|n_{y}-\tilde{n}_{y}\right|}{N_{y}-1}W_{y}\right)^{2}}\right),\end{array}$ (1)
where $j_{0}(\cdot)$ denotes the spherical Bessel function of the first kind, which can be approximated by $\mathrm{sinc}(x)={\sin x}/{x}$, and $\lambda$ represents the carrier wavelength. The downlink channel for a typical user can then be modeled as
$\mathbf{g}=\delta\mathbf{U}_{s}\sqrt{\mathbf{\Lambda}_{s}}\mathbf{G}+\mathbf{n},$ (2)
where $\mathbf{U}_{s}\in\mathbb{C}^{N_{s}\times N_{s}}$ and $\mathbf{\Lambda}_{s}\in\mathbb{C}^{N_{s}\times N_{s}}$ are defined in accordance with [20], representing the eigenvector matrix and eigenvalue diagonal matrix obtained from the eigen-decomposition of the spatial correlation matrix $\mathbf{J}_{s}=\mathbf{U}_{s}\mathbf{\Lambda}_{s}\mathbf{U}_{s}^{H}$ , whose entries are given by (1). The matrix $\mathbf{G}\in\mathbb{C}^{N_{s}\times M_{t}}$consists of independent complex Gaussian random variables with zero mean and variance of $1/2$. The parameter $\delta$ characterizes the path loss, and ${\bf n}$ is the additive white Gaussian noise.
III Proposed Approach
Here, we propose CANet, a noise-robust architecture for the channel extrapolation task. We design a CAB to locally fuse the features of known CSI with weighted integration, while constraining the propagation of features from unknown CSI. To enhance global modeling capability, we introduce a CSCA module to facilitate feature completion for large unobserved port regions. Moreover, we adopt ConvNeXt v2 with dropout regularization to improve feature extraction and generalization after feature completion [31]. A spatial amplitude perturbation module is also incorporated to enhance the model’s adaptability and robustness under various noise conditions.
III-A Masking Strategy
We denote the perfect CSI as $\mathbf{g}\in\mathbb{C}^{N_{s}\times M_{t}}$, and construct a tensor $\mathbf{u}\in\mathbb{R}^{2M_{t}\times N_{y}\times N_{x}}$ to represent the concatenated real and imaginary components of the CSI across a 2D fluid antenna grid of size $N_{x}\times N_{y}$. To simulate partial observations, a subset of antenna ports is randomly masked along the spatial grid, with a masking ratio typically ranging from $80\%$ to $95\%$. The resulting masked tensor is denoted as $\tilde{\mathbf{u}}\in\mathbb{R}^{2M_{t}\times N_{y}\times N_{x}}$, in which the unobserved positions are replaced with a distinct negative value (e.g., $-10$) to make them distinguishable for the network. Additionally, we define a binary indicator matrix $\mathbf{E}_{\text{flag}}\in\{0,1\}^{N_{y}\times N_{x}}$, where a value of $1$ indicates an observed port and 00 is a masked (unobserved) port. No prior knowledge of the channel is used to select ports for extrapolation. Instead, a random masking strategy is employed to ensure flexibility and enhance the generalization capability of the model.
Figure 1: The proposed CANet for CSI extrapolation in FAS channels, where $\oplus$, $\otimes$ and $\odot$ denote fuse, dot product and element-wise product, respectively.
III-B Model Architecture
The overall architecture of CANet consists of four training modules. Next, we summarize the role of each module.
III-B1 CAB
For channel extrapolation tasks with strong spatial correlations, using convolutional neural network is a natural choice for feature extraction. However, conventional convolution is susceptible to interference from unobserved ports within the receptive field, which can degrade the quality of feature representations. To overcome this issue, we propose a dual-branch architecture that separately models and predicts the CSI of observed and unobserved ports, thereby enhancing both the modeling accuracy and the robustness of extrapolation.
Specifically, we first compute a soft mask (confidence map) over the entire feature map using convolution. This soft mask reflects the reliability of each feature. [^1] Features affected by masked ports are assigned lower confidence, while those from observed regions receive higher confidence. Then we perform a weighted fusion of the outputs from the convolutional and attention branches to further refine feature extraction within the receptive field. For each pixel $x_{o}\in\mathbb{R}^{C}$ on the feature map $F\in\mathbb{R}^{C\times H\times W}$, its corresponding soft mask value $m$ is computed by the following convolutional operation:
$m=\sigma\left(\sum_{i\in\mathcal{B}_{o}}w_{i}^{m}x_{i}\right),$ (3)
where $\sigma$ denotes the sigmoid activation function, $x_{i}$ represents the input features within the receptive field, $w_{i}$ denotes learnable convolutional filters, and $\mathcal{B}_{o}$ represents the receptive field range corresponding to filter $w^{m}$. The magnitude of $m$ signifies the confidence level of the feature at that location. Conditioned on the learned soft mask $m$, we then introduce a self-attention mechanism [32] to derive a more contextually relevant feature representation $x^{a}$, given by
$x^{a}=\sum_{i\in\mathcal{B}_{o}}\left\langle mw^{q}x_{o},mw^{k}x_{i}\right\rangle mx_{i}w^{v}x_{i},$ (4)
[…]
$m\in\sum_{i\in\mathcal{B}_{o}}w_{i}^{c}x_{i}.$ (5)
Finally, we perform a weighted fusion of the convolutional branch output $x_{c}$ and the attention branch output $x_{a}$ based on the soft mask $m$, yielding the output
$o=m\cdot x^{c}+(1-m)\cdot x^{a}.$ (6)
III-B2 CSCA
In the CSCA module, we first reduce the input feature channels to half of the original before feeding them into two self-attention modules. One branch directly undergoes self-attention at the same scale, while the other duplicates the feature map and down-samples it to half of the original resolution. The downsampled feature map is then processed along with the original-scale features within the self-attention mechanism. At each self-attention stage, patches are extracted from the input features, and the cosine similarity between patches in the missing region and those in the external region is computed to establish cross-region feature correlations. Formally, let $q_{i}$ denote the $i$-th patch from the observed region $\psi^{l}$, and $k_{j}$ denote the $j$-th patch from the masked region $\phi^{l}$. The cosine similarity is computed as
$s_{i,j}=\left\langle\frac{q_{i}}{\|q_{i}\|},\frac{k_{j}}{\|k_{j}\|}\right\rangle.$ (7)
The attention weights are then computed by
$a_{i,j}=\frac{\exp(s_{i,j})}{\sum_{j}\exp(s_{i,j})}.$ (8)
The unobserved port patches are then reconstructed based on the computed attention scores, given by
$\hat{q}=\sum_{j}a_{i,j}v_{j},$ (9)
where $v_{j}$ represents the $j$-th patch extracted from features outside unobserved regions. Finally we concatenate the results of the two branches as the final result.
TABLE I: Architecture of the proposed CANet
Layer Operator Input Size Output Size Output Channel Kernel Size Stride Nonlinearity
1 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $64$ $5$ $1$ LeakyReLU
2 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $2$ LeakyReLU
3 CAB $32~{}\times~{}6$ $16~{}\times~{}8$ $128$ $3$ $1$ LeakyReLU
4 CAB $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ $3$ $1$ LeakyReLU
5 CAB $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ $3$ $1$ LeakyReLU
6 CSCA $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ Null Null Null
7–16 ConvNeXt v2 $16~{}\times~{}8$ $16~{}\times~{}8$ $256$ $7$ $3$ GeLU
17 BilinerUp $16~{}\times~{}8$ $32~{}\times~{}16$ $256$ $3$ $1$ Null
18 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $1$ LeakyReLU
19 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $128$ $3$ $1$ LeakyReLU
20 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $64$ $3$ $1$ LeakyReLU
21 CAB $32~{}\times~{}16$ $32~{}\times~{}16$ $64$ $3$ $1$ LeakyReLU
22 Conv $32~{}\times~{}16$ $32~{}\times~{}16$ $16$ $7$ $1$ Null
III-B3 ConvNeXt v2 with dropout regularization
After CAB and CSCA, we adopt ConvNeXt V2 [31] with dropout regularization as the backbone network to further enhance feature extraction and improve generalization. ConvNeXt has been widely used for its strong feature representation capabilities, and ConvNeXt V2 builds upon this by introducing the global response normalization (GRN) module, which enables more expressive feature representations. In our design, we employ the ConvNeXt V2 block for channel extrapolation. By randomly dropping a subset of neurons during training, dropout alters the effective network structure in each training epoch, leading to a degree of statistical independence across different training instances, thereby reducing prediction variance and enhancing generalization under diverse noise conditions.
III-B4 Spatial amplitude perturbation
To enhance the model’s robustness, we propose a spatial amplitude perturbation strategy at the input stage, inspired by frequency perturbation techniques in the image domain [33, 34]. Specifically, the input CSI features are first zero-mean normalized to remove the DC component, then transformed into the frequency domain via FFT. The amplitude spectrum is then perturbed in a stochastic manner based on a predefined probability, effectively simulating potential frequency interference and noise. This approach improves the model’s adaptability to unseen distributions:
$\tilde{\mathbf{z}}=\tilde{\mathbf{u}}-\frac{1}{HW}\sum_{h=1}^{H}\sum_{w=1}^{W}\tilde{\mathbf{u}}_{h,w}.$ (10)
$\hat{\mathbf{z}}=\mathcal{F}^{-1}\left(\mathcal{F}(\tilde{\mathbf{z}})\cdot\left(1+\bm{\gamma}\cdot\bm{\xi}\cdot\bm{\delta}\right)\right)+\bar{\mathbf{z}},$ (11)
where $\tilde{\mathbf{z}}$ is the spatially zero-mean feature, $\hat{\mathbf{z}}$ is the final perturbed feature after reconstruction, $\mathcal{F}(\cdot)$ and $\mathcal{F}^{-1}(\cdot)$ denote the 2D discrete Fourier transform and its inverse. The perturbation mask $\bm{\delta}\in\{0,1\}^{H\times W}$ above is sampled from a Bernoulli distribution with probability $\mu$, and the amplitude coefficients $\bm{\xi}\in[0,1]^{H\times W}$ are sampled from a normal distribution $\mathcal{N}(0,1)$. The hyperparameters $\mu$ and $\gamma$ control the perturbation probability and maximum strength, respectively. Only the amplitude of each frequency component is perturbed, while the phase is fixed to maintain structural consistency.
III-B5 Loss Function
Since the CSI tensor is represented by concatenating the real and imaginary parts, we reconstruct it into a complex-valued tensor before computing the frequency-domain loss to enable the 2D Fourier transform. We denote the original CSI tensor as $\mathbf{{U}_{c}}\in\mathbb{R}^{M_{t}\times N_{y}\times N_{x}}$, and $\hat{\mathbf{{U}_{c}}}\in\mathbb{R}^{M_{t}\times N_{y}\times N_{x}}$ denote the predicted CSI recovered from the masked input $\tilde{\mathbf{{U}_{c}}}$. Let $\Omega\subset\{1,\dots,N_{y}\}\times\{1,\dots,N_{x}\}$ denote the set of unobserved spatial locations, with $|\Omega|=N_{a}$. The total loss is then defined as
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{MSE}}+\beta\mathcal{L}_{\mathrm{FFT}},$ (12)
where $\mathcal{L}_{\text{MSE}}$ and $\mathcal{L}_{\mathrm{FFT}}$ are the mean-square-error (MSE) loss and the FFT loss attributed to spatial amplitude perturbation, respectively, calculated as
$\mathcal{L}_{\text{MSE}}$ $=\mathbb{E}\left[\frac{1}{|\Omega|}\sum_{(y,x)\in\Omega}\left\|\hat{\mathbf{{U}_{c}}}[:,y,x]-\mathbf{{U}_{c}}[:,y,x]\right\|_{2}^{2}\right],$ (13)
$\mathcal{L}_{\mathrm{FFT}}$ $=\mathbb{E}\left[\left\|\left|\mathcal{F}(\hat{\mathbf{{U}_{c}}})\right|-\left|\mathcal{F}(\mathbf{{U}_{c}})\right|\right\|_{2}^{2}\right],$ (14)
in which $\beta$ balances the contribution of the two losses. The FFT loss enforces global consistency in the spatial frequency domain, which helps preserve large-scale channel structure.
IV Numerical Results
In the simulations, the FAS channel involves $U_{t}$ users. Each user has a planar FAS of size $W_{s}\,~{}\text{cm}^{2}$ and $N_{s}$ antenna ports. We assume $m_{t}=U_{t}$, and the BS is equipped with $M_{t}$ FPAs. The carrier frequency is $3.4~{}{\rm GHz}$. Detailed simulation parameters are summarized in Table I. The dataset used for analysis contains a total of $310,000$ samples, divided into $240,000$ training samples, $60,000$ validation samples, and $10,000$ test samples for model training and performance evaluation. The total number of ports at the FAS side is set to $N_{s}=512$.
The performance is evaluated through the normalized MSE (NMSE), defined as
$\text{NMSE}(\hat{\mathbf{g}},\mathbf{g})=\frac{\sum_{n=1}^{N_{\text{test}}}\left\|\mathbf{g}_{n}-\hat{\mathbf{g}}_{n}\right\|^{2}}{\sum_{n=1}^{N_{\text{test}}}\left\|\mathbf{g}_{n}\right\|^{2}},$ (15)
where $N_{\text{test}}$ denotes the number of test samples. The predicted CSI matrix $\hat{\mathbf{g}}$ is obtained by rearranging the extrapolated tensor $\hat{\mathbf{U}}_{c}$ into the standard CSI format. The batch size used in training is set to $32$. The initial learning rate is $0.0005$. For optimization, the AdamW optimizer is employed with the parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and the weight decay set to $0.0001$. We set the maximum strength $\gamma$ to $0.5$, $\mu$ to $0.05$, and the dropout rate to $0.5$, $\beta$ to $0.02$.
(a) $W_{s}=2~{}{\rm cm}\times 4~{}{\rm cm}$
(b) $W_{s}=8~{}{\rm cm}\times 16~{}{\rm cm}$
Figure 2: NMSE versus the number of observed ports with different FAS sizes.
Figs. 2(a) and 2(b) illustrate the NMSE curves of different models under various SNR conditions, as a function of the observed port ratio, for FAS sizes of $2~{}{\rm cm}\times 4~{}{\rm cm}$ and $8~{}{\rm cm}\times 16~{}{\rm cm}$, respectively. We reproduce the AGMAE in [25] as a baseline for comparison. The simplified version of CANet without spatial amplitude perturbation and spectral consistency loss is denoted as ‘CANet-B’, while the full CANet model includes all the proposed components. In the figures, curves with the same line style represent the same SNR setting.
The results demonstrate that as the ratio of observed ports increases, the NMSE of all the models decreases considerably, showing that richer observed information leads to better reconstruction performance. Under extremely low observation rates (e.g., $5\%$), AGMAE slightly outperforms CANet-B, suggesting that its structure has certain advantages under highly sparse inputs. But in most observation rates, CANet-B consistently surpasses AGMAE. Furthermore, with the incorporation of amplitude spectrum perturbation and spectral consistency loss, the overall performance of CANet is further improved. For example, at an observation rate of $20\%$ with the $2~{}{\rm cm}\times 4~{}{\rm cm}$ FAS, CANet achieves up to a $3~{}{\rm dB}$ NMSE gain over CANet-B, demonstrating stronger noise robustness and generalization ability. Fig. 2(b) shows a similar trend under a larger FAS size, validating the adaptability and effectiveness of the proposed method across different physical FAS configurations.
V Conclusions
In this correspondence, we proposed CANet, a robust CSI extrapolation model that excels in processing unobserved port positions in FAS. CANet integrates a CAB, CSCA, and a ConvNeXt v2 backbone. This combination allows for adaptive CSI feature extraction at both local and global levels. A spatial amplitude perturbation strategy was also proposed to enhance the loss function with an FFT loss. Our simulation results indicated that CANet consistently outperforms benchmark models, demonstrating improved accuracy and remarkable robustness in channel extrapolation under various noise conditions.